OpenStudy (anonymous):

4. a. Calculate the scalar projection of (5,3) onto (7,-2).

OpenStudy (anonymous):

\[=(5*7-3*2)/\sqrt{53}\]

OpenStudy (anonymous):

can you explain that? i dont understand. @myko

OpenStudy (anonymous):

To calculate the projection of a onto b, 1) Do the dot product. 2) Divide by the length of b.

OpenStudy (anonymous):

Do you know how to do the dot product?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

<a,b><c,d>=ac+bd this is scalar product of two vectors

OpenStudy (anonymous):

It's pretty simple =) If I have \[\left(\begin{matrix}a \\ b\end{matrix}\right) . \left(\begin{matrix}c \\ d\end{matrix}\right)\] I do a*c + b*d

OpenStudy (amistre64):

im more concerned that your asking this question without knowing how to do the fundamentals of it :/

OpenStudy (anonymous):

I'm not concerned. She's asking for an explanation and sticking around to hear it.

OpenStudy (anonymous):

good one

OpenStudy (anonymous):

Does that make sense, sea-child?

OpenStudy (anonymous):

Could you try to do this dot product for me? \[\left(\begin{matrix}2 \\ -3\end{matrix}\right).\left(\begin{matrix}4 \\ -6\end{matrix}\right)\]

OpenStudy (anonymous):

i got 26.

OpenStudy (anonymous):

i dont know if thats right

OpenStudy (anonymous):

Yes, perfect =)

OpenStudy (amistre64):